OpenStudy (nurali):

A die is a cube with the numbers 1, 2, 3, 4, 5, and 6 written on its six faces,one number per face. A die is rolled 12 times. Here are the faces that show: 6, 3, 1, 6, 3, 5, 3, 1, 2, 1, 3, 2. Which face was the mode?

OpenStudy (phi):

do you know the definition of mode ?

OpenStudy (phi):

The mode in a list of numbers refers to the number that occurs most frequently.

OpenStudy (phi):

The mode is not necessarily unique

OpenStudy (phi):

remember Mode is MOST

OpenStudy (phi):

if there is a tie for most, then all the numbers in the tie can be called the mode.

OpenStudy (phi):

one way is sort the list. I think 3 is the only mode (if I counted correctly)

hartnn (hartnn):

3 repeats 5 times

OpenStudy (phi):

there are 5 3's so it is the mode (3 occurs most often in the data)

hartnn (hartnn):

mode = 3, number of modes =1

hartnn (hartnn):

number of modes was asked, so answer =1

hartnn (hartnn):

add all them, thats sum count how many numbers are there, that 'n' average = sum/n

hartnn (hartnn):

let me calculate

hartnn (hartnn):

yes, thats correct.

OpenStudy (nurali):

The stem and leaf plot below shows the distribution of weights (in ounces) of newborns at a hospital. The line 10 | 358 should be read as 103, 105, 108. Select the option that is closest to the average of these weights. Yes, you can enter the list into your calculator and compute the average. But it will be a lot faster to do this exercise by inspection. 9 | 2 10 | 358 11 | 034445559 12 | 002258 13 | 12678 14 | 013446

hartnn (hartnn):

there are so many weights!

hartnn (hartnn):

what are your choices ?

hartnn (hartnn):

idk how to do it with inspection..with calculator, it'll take many minutes!

hartnn (hartnn):

i can guess it to be 125

hartnn (hartnn):

i can't be sure because i don't know how to do it with inspection, but since there were many numbers with "11. ",the average has to be near 110 or greater, so i guessed 125

OpenStudy (nurali):

A surveyor measures the height of a hill 6 times and gets an average of 268 meters. Previously, he had measured the height 10 times and got an average measurement of 268.5 meters. Find the average (in meters) of all 16 measurements.

hartnn (hartnn):

(268*6+268.5*10)/(10+6)

hartnn (hartnn):

is than an option ? if yes, thats correct.

OpenStudy (nurali):

"Household size" is defined as the number of people living in a household. In a sample of 1200 households in California in 2010, the average household size was 2.9 people. Which of the following conclusions can be drawn from this information? Select all that are correct.

OpenStudy (nurali):

a) There must have been an error in the calculation, because no household can have 2.9 people. b) The typical household in the sample consisted of 2.9 people. c) There were 3480 people living in the sampled households. d) In the sample, 600 households had fewer than 2.9 people.

hartnn (hartnn):

c) is true for sure.

hartnn (hartnn):

a) and b) are false, nothing can be said for d)

OpenStudy (nurali):

As part of a fund raising campaign, the 12 members of the governing board of an organization pledge to raise $100,000. In the first month of the campaign, the average amount raised by the 12 board members is $5,200. How much more must they raise, on average, so that they can keep their pledge? [Answer format: Your answer will be interpreted as dollars and cents. If you think the answer is $12.34, enter 12.34]

hartnn (hartnn):

total raised = 12*5200 remaining = 100000-12*5200

hartnn (hartnn):

average = remaining/12

hartnn (hartnn):

remaining = 100000-12*5200=37600 average = 37600/12 = 3133.34 is this an option ?

hartnn (hartnn):

i am done for today, good luck with rest of the questions :) ....

OpenStudy (nurali):

help please.

OpenStudy (nurali):

@campbell_st

OpenStudy (campbell_st):

try posting each question separately... makes it easier for people to help you.